# Revisiting the Concordance $\Lambda$CDM model using Gamma-Ray Bursts together with Supernovae Ia and Planck data

### v1

@@ Abstract
The Hubble constant, $H_{0}$, tension is the tension among the local probes, Supernovae Ia, and the Cosmic Microwave Background Radiation. It has been almost a decade, and this tension still puzzles the community. Here, we add intermediate redshift probes, such as Gamma-Ray Bursts (GRB) and Quasars (QS0s), to check if and to what extent these higher redshift probes can reduce this tension. We use the three-dimensional fundamental plane relation among the prompt peak luminosity, the luminosity at the end of the plateau emission, and its rest frame duration. We find similar trend in GRB intrinsic parameters as previously seen in Pantheon-Plus intrinsic parameters. We find an apparent $3.14\sigma$ tension for the GRB intrinsic parameter $b$. Indeed, this tension disappears and the parameters are actually compatible within $2.26\sigma$. Another interesting point is that the 3D relation plays an important role in conjunction with Supernovae data with Pantheon Plus and that this apparent discrepancy show how it is important the correction for selection biases and redshift evolution. The incorporation of redshift evolution correction results in a reduction of the GRB tension to $2.26\sigma$ when adjusting correction parameters. We envision that with more data this indication of tension will possibly disappear when the evolutionary parameters of GRBs are computed with increased precision.
@@ 1 Introduction
The standard concordance $\Lambda$CDM model constituting the Dark Energy term as a cosmological constant (CC) $\Lambda$ with the cold dark matter has been a pivotal framework in explaining both the early time observations [1, 2] as well as late time observations [3, 4] providing a solid foundation for understanding the observed acceleration of the Universe. Despite its success, the intrinsic nature of the dark energy component in this model remains a mystery. It is commonly characterized by negative pressure and positive energy density, a framework consistent with the requirements for cosmic acceleration [5]. The $\Lambda$ constant has been a fitting candidate, arising from the vacuum energy density of the Universe rooted in quantum field theory [6]. However, the theoretical construction needs to be revised in accounting for the observed value, introducing a substantial discrepancy of the order of $\mathcal{O}(121)$ [7]. Despite these theoretical challenges, the $\Lambda$CDM model stands as the most favoured choice in accommodating a diverse range of cosmological observations. Nevertheless, recent years have witnessed emerging inconsistencies within the model, particularly in measuring various cosmological parameters from different observational scenarios. Notable examples include tensions affecting the Hubble constant $H_{0}$ [8, 9, 10, 11, 12, 13, 14, 15, 16, 17] and the amplitude of matter fluctuations $S_{8}$ [18, 19, 20]. These discrepancies suggest the possibility of new physics being at play within the Dark Energy sector, calling for a more comprehensive exploration of alternative models and theoretical frameworks to pursue a more accurate cosmological description.
@@ 2.1 Background
where $H_{0}$ is the Hubble constant, and $\Omega_{r},\Omega_{m},\Omega_{\Lambda}$ are the present day values for the density parameters of radiation, matter and dark energy respectively. In our work we have considered the Universe to be flat as precisely predicted by Planck experiments, which gives the constraint on the density parameters as follows:
@@ 2.2 Data and the corresponding likelihoods
  • CMB: From the Planck 2018 legacy data release, we use the CMB measurements, namely the $-\ell$ Plik TT likelihood (in the multipole range $30\leq\ell\leq 2508$), TE and EE (in the multipole range $30\leq\ell\leq 1996$), $-\ell$ TT-only ($2\leq\ell\leq 29$), the $-\ell$ EE-only ($2\leq\ell\leq 29$) likelihood [57], in addition to the CMB lensing power spectrum measurements [58]. We refer to this dataset as Planck. We used the methodology detailed in [59] to compute the likelihood of the CMB power spectrum.
@@
  • Pantheon-Plus: We use the SNe Ia distance moduli measurements from the Pantheon-Plus sample [60], which consists of 1550 distinct SNe Ia ranging in the redshift interval $z\in[0.001,2.26]$ and in total has 1701 with duplicated entries. We refer to this dataset as Pantheon-Plus. We also consider the SH0ES Cepheid host distance anchors, which facilitate constraints on both $M$ and $H_{0}$. When utilizing SH0ES Cepheid host distances, the SNe Ia distance residuals are modified following the relationship eq. (14) of [60]. We refer to this dataset as Pantheon-Plus&SH0ES. The likelihood of Pantheon-Plus is based on the measurement of distance modulus ($\mu$), which is given by Tripp’s formula [61] as:
@@
    $-2{\rm ln}(\mathcal{L})=\chi^{2}=\Delta\vec{D}^{T}~{}C_{\rm stat+syst}^{-1}~{}\Delta\vec{D},$ (2.4)
@@
    $-2{\rm ln}(\mathcal{L}^{\prime})=\Delta\vec{D^{\prime}}^{T}~{}(C^{\rm SN}_{\rm stat+syst}+C_{\rm stat+syst}^{\rm Cepheid})^{-1}~{}\Delta\vec{D^{\prime}},$ (2.9)
@@
    The creation of the Platinum Sample involves additional requirements for the plateau: 1) the end time of the plateau ($T_{X}$) should not coincide with observational gaps in the light curves (LCs), ensuring direct determination from data rather than relying on LC fitting; 2) the plateau duration must be at least 500 seconds so that we do not include dubious cases where the plateau is masked out; and 3) no flares should occur at the onset or throughout the entire plateau duration. This additional criteria refines the criteria introduced by [27].
    Applying these stringent criteria results in a Platinum Sample consisting of 50 gamma-ray bursts (GRBs) out of the 222 plateaus analyzed, with the farthest GRB in this sample located at redshift $z=5$. The luminosities both in the prompt and in the plateaus and the times at the end of the plateaus carry error bars which are comparable, namely the $\frac{\Delta_{x}}{x}$, $\frac{\Delta_{y}}{y}$ and $\frac{\Delta_{z}}{z}$ are of the same order of magnitude, where $\Delta_{x}$ is the error on the x-axis (error on time in our case) and $\Delta_{y}$ is the error on y axis (luminosity at the end of the plateau phase) and $\Delta_{z}$ is the uncertainty on the z-axis, which is the peak luminosity. Thus, it is necessary to adopt methods which take into account both error bars like the [66, 67]. In addition, this method allows us to introduce an intrinsic scatter, $\sigma_{int}$. The fundamental plane relation has the following form:
    $\log L_{X}=c+a\cdot\log T^{*}_{X}+b\cdot(\log L_{peak}),$ (2.10)
    We compare both $\mu_{obs,GRBs}$ with the theoretical $\mu_{th}$, defined as:
@@
    By combining the definitions in Equ. 2.10 and Equ. 2.12 we finally reproduce Equation 2.11. We write as follows the likelihoods:
@@
    $\mathrm{log}L_{X}=g\,\mathrm{log}L_{UV}+b$ (2.15)
    where $L_{X}$ and $L_{UV}$ are the luminosities (in $\mathrm{erg\,s^{-1}\,Hz^{-1}}$) at 2 keV and 2500 Å, respectively. In [75] this relation has been corrected for selection biases and redshift evolution using the EP and it has been proved that it is intrinsic to QSO’s properties so that it can be reliably used to standardize QSOs as cosmological tools. To make use of Equation (2.15), we compute luminosities from measured flux densities $F$ according to $L_{X,UV}=4\pi d_{L}^{2}\ F_{X,UV}$, where $d_{L}$ is computed under the assumption of a cosmological model using the corresponding parameters as free parameters of the fit. Usually, luminosities of the sources have to be corrected for the K-correction, defined as $1/(1+z)^{1-\beta}$, where $\beta$ is the spectral index of the sources when we assume a simple power law for the spectrum, but for QSOs it is assumed to be $\beta=1$, leading to a $K=1$. So the K-correction has been omitted following [70]. The likelihood function (LF) used for QSOs is [see also 76, 77, 78, 79]:
    $\text{ln}(LF)_{\text{QSO}}=-\frac{1}{2}\sum_{i=1}^{N}\left[\frac{(y_{i}-\phi_{i})^{2}}{s^{2}_{i}}+\text{ln}(s^{2}_{i})\right].$ (2.16)
@@
Along with the 6 standard parameters in baseline $\Lambda$CDM model we have $1$ intrinsic parameter for Pantheon-Plus i.e. $M_{B}$, 3 intrinsic parameters for GRB i.e. $a,b,svGRB$ and 3 intrinsic parameters for QSO i.e. $g,be,svQ$. We ran our modified version of the publicly available CLASS+MontePython code [80, 81, 82] that implements the model described in 2.1, using Metropolis-Hastings algorithm to derive constraints on cosmological parameters as well as the intrinsic parameters. All of our runs were set up to reach a Gelman-Rubin convergence criterion of $R-1<10^{-2}$.
@@ 3.1 $M$ Tension
Figure 1: Contour plots of the values of $M$ and $H_{0}$ using the Planck+ Pantheon-Plus data and Pantheon-Plus combined with SHOES are shown in dark green and light green, respectively.
@@ 3.3 Impact on GRB intrinsic parameters
Figure 3: The contour plots for the parameters $a$, $b$, $\sigma_{v}$, and $H_{0}$ for Planck combined with GRBs and for GRBs combined with Pantheon and with SHOES data are shown in darker blue and lighter green, respectively.

### v2

@@ Abstract
The Hubble constant, $H_{0}$, tension is the tension among the local probes, Supernovae Ia, and the Cosmic Microwave Background Radiation. This tension has persisted for decades and continues to puzzle the community. Here, we add intermediate redshift probes, such as Gamma-Ray Bursts (GRB) and Quasars (QS0s), to check if and to what extent these higher redshift probes can reduce this tension. We use the three-dimensional fundamental plane relation among the prompt peak luminosity, the luminosity at the end of the plateau emission, and its rest frame duration. We find similar trend in GRB intrinsic parameters as previously seen in Pantheon-Plus intrinsic parameters. We find an apparent $3.14\sigma$ tension for the GRB intrinsic parameter $b$. Indeed, this tension disappears and the parameters are actually compatible within $2.26\sigma$. Another interesting point is that the 3D relation plays an important role in conjunction with Supernovae data with Pantheon Plus and that this apparent discrepancy shows the importance of the correction for selection biases and redshift evolution. The incorporation of redshift evolution correction results in a reduction of the GRB tension to $2.26\sigma$ when adjusting correction parameters. We envision that with more data this indication of tension will possibly disappear when the evolutionary parameters of GRBs are computed with increased precision.
@@ 1 Introduction
The standard concordance $\Lambda$CDM model constituting the Dark Energy term as a cosmological constant (CC) $\Lambda$ with the cold dark matter has been a pivotal framework in explaining both the early time observations [1, 2] as well as late time observations [3, 4] providing a solid foundation for understanding the observed acceleration of the Universe. Despite its success, the intrinsic nature of the dark energy component in this model remains a mystery. It is commonly characterized by negative pressure and positive energy density, a framework consistent with the requirements for cosmic acceleration [5]. The $\Lambda$ constant has been a fitting candidate, arising from the vacuum energy density of the Universe rooted in quantum field theory [6]. However, the theoretical construction needs to be revised in accounting for the observed value, introducing a substantial discrepancy of the order of $\mathcal{O}(121)$ [7]. Despite these theoretical challenges, the $\Lambda$CDM model stands as the most favoured choice in accommodating a diverse range of cosmological observations. Nevertheless, recent years have witnessed emerging inconsistencies within the model, particularly in measuring various cosmological parameters from different observational scenarios. Notable examples include tensions affecting the Hubble constant $H_{0}$ [8, 9, 10, 11, 12, 13, 14, 15, 16, 17] and the amplitude of matter fluctuations $S_{8}$ [18, 19, 20]. These discrepancies suggest the possibility of new physics being at play within the Dark Energy sector, calling for a more comprehensive exploration of alternative models and theoretical frameworks to pursue a more accurate cosmological description.[Eleonora2021, Kamionkowski2022, Sunny2023, Sen:2021wld]
@@ 2.1 Background
where $H_{0}$ is the Hubble constant, and $\Omega_{r},\Omega_{m},\Omega_{\Lambda}$ are the present day values for the density parameters of radiation, matter and dark energy respectively. In our work we have considered the Universe to be flat as precisely predicted by Planck experiments [1], which gives the constraint on the density parameters as follows:
@@ 2.2 Data
  • CMB: From the Planck 2018 legacy data release, we use the CMB measurements, namely the $-\ell$ Plik TT likelihood (in the multipole range $30\leq\ell\leq 2508$), TE and EE (in the multipole range $30\leq\ell\leq 1996$), $-\ell$ TT-only ($2\leq\ell\leq 29$), the $-\ell$ EE-only ($2\leq\ell\leq 29$) likelihood [57], in addition to the CMB lensing power spectrum measurements [58]. We refer to this dataset as Planck.
@@
  • Pantheon-Plus: We use the SNe Ia distance moduli measurements from the Pantheon-Plus sample [60], which consists of 1550 distinct SNe Ia ranging in the redshift interval $z\in[0.001,2.26]$ and in total has 1701 with duplicated entries. We refer to this dataset as Pantheon-Plus. We also consider the SH0ES Cepheid host distance anchors, which facilitate constraints on both $M$ and $H_{0}$. When utilizing SH0ES Cepheid host distances, the SNe Ia distance residuals are modified following the relationship eq. (14) of [60]. We refer to this dataset as Pantheon-Plus&SH0ES.
    The creation of the Platinum Sample involves additional requirements for the plateau: 1) the end time of the plateau ($T_{X}$) should not coincide with observational gaps in the light curves (LCs), ensuring direct determination from data rather than relying on LC fitting; 2) the plateau duration must be at least 500 seconds so that we do not include dubious cases where the plateau is masked out; and 3) no flares should occur at the onset or throughout the entire plateau duration. These additional criteria refines the criteria introduced by [27].
    Applying these stringent criteria mentioned above, we obtain a Platinum Sample consisting of 50 GRBs out of the 222 X-ray plateaus analyzed. The farthest GRB in this sample is located at redshift $z=5$. The luminosities both in the prompt and in the plateaus and the times at the end of the plateaus carry error bars which are comparable, namely the $\frac{\Delta_{x}}{x}$, $\frac{\Delta_{y}}{y}$ and $\frac{\Delta_{z}}{z}$ are of the same order of magnitude, where $\Delta_{x}$ is the error on the x-axis (error on time in our case) and $\Delta_{y}$ is the error on y axis (luminosity at the end of the plateau phase) and $\Delta_{z}$ is the uncertainty on the z-axis, which is the peak luminosity. Thus, it is necessary to adopt methods which take into account both error bars like the [66, 67]. In addition, this method allows us to introduce an intrinsic scatter, $\sigma_{int}$. The fundamental plane relation has the following form:
    $\log L_{X}=c+a\cdot\log T^{*}_{X}+b\cdot(\log L_{peak}),$ (2.3)
    $\mathrm{log}L_{X}=g\,\mathrm{log}L_{UV}+b$ (2.6)
    where $L_{X}$ and $L_{UV}$ are the luminosities (in $\mathrm{erg\,s^{-1}\,Hz^{-1}}$) at 2 keV and 2500 Å, respectively. In [75] this relation has been corrected for selection biases and redshift evolution using the EP and it has been proved that it is intrinsic to QSO’s properties so that it can be reliably used to standardize QSOs as cosmological tools. To make use of Equation (2.6), we compute luminosities from measured flux densities $F$ according to $L_{X,UV}=4\pi d_{L}^{2}\ F_{X,UV}$, where $d_{L}$ is computed under the assumption of a cosmological model using the corresponding parameters as free parameters of the fit. Usually, luminosities of the sources have to be corrected for the K-correction, defined as $1/(1+z)^{1-\beta}$, where $\beta$ is the spectral index of the sources when we assume a simple power law for the spectrum, but for QSOs it is assumed to be $\beta=1$, leading to a $K=1$. So the K-correction has been omitted following [70].
2.3 Methodology
We have performed the statistical analysis using the Bayesian inference technique, which is extensively used for parameter estimation in cosmological models. Bayesian inference provides a powerful framework for updating the probability of a hypothesis as more evidence or information becomes available. In this context, the posterior probability distribution function of the model parameters is determined by combining the likelihood function with the prior probability of the model parameters. Mathematically, the posterior distribution is proportional to the product of the likelihood function and the prior distribution. To estimate the model parameters, we commonly use multivariate Gaussian likelihoods. These likelihoods quantify how well the model parameters explain the observed data and are given by the expression:
${\cal L}(\Theta)\propto\exp\left[-\frac{\chi^{2}(\Theta)}{2}\right],$ (2.7)
where $\Theta$ represents the set of model parameters. Here, $\chi^{2}(\Theta)$ is the chi-squared statistic, which measures the difference between the observed data and the values predicted by the model for a given set of parameters. In our analysis, we have adopted a uniform prior distribution for the model parameters. A uniform prior indicates that all parameter values are initially considered equally likely before observing the data. Under this assumption, the posterior probability distribution simplifies to being directly proportional to $\exp\left[-\frac{\chi^{2}(\Theta)}{2}\right]$. Therefore, minimizing the chi-squared statistic $\chi^{2}(\Theta)$ maximizes the likelihood function and, consequently, the posterior probability. This means that the parameter set $\Theta$ that results in the smallest $\chi^{2}(\Theta)$ is the most probable given the observed data. The likelihoods for all corresponding data sets are defined in a similar manner, ensuring that we can comprehensively evaluate the goodness-of-fit for different parameter sets across various observations. We have defined the different likelihoods as the following:
  • Planck: The Planck likelihood is associated with the CMB temperature and polarization spectra. It is defined in two multipole ranges: $\ell\leq 30$ and $\ell\geq 30$. For $-\ell$ values, the likelihood has two components: 1) the baseline Plik likelihood and 2) the CamSpec likelihood. For $-\ell$ values, it consists of the $-\ell$ temperature Commander likelihood and the $-\ell$ SimAll polarization likelihood. Additionally, we have utilized the lensing likelihood of the CMB spectrum. We used the publicly available Planck likelihood code [^1] Planck Likelihood code to calculate these likelihoods. Detailed implementation can be found in Section 2.2 in [1]; see also [57].
  • Pantheon-Plus:The likelihood of Pantheon-Plus is based on the measurement of distance modulus ($\mu$), which is given by Tripp’s formula [61] as:
@@ 2.3 Methodology
    $-2{\rm ln}(\mathcal{L}(SN))=\chi^{2}=\Delta\vec{D}^{T}~{}C_{\rm stat+syst}^{-1}~{}\Delta\vec{D},$ (2.9)
@@
    $-2{\rm ln}(\mathcal{L}(SN)^{\prime})=\Delta\vec{D^{\prime}}^{T}~{}(C^{\rm SN}_{\rm stat+syst}+C_{\rm stat+syst}^{\rm Cepheid})^{-1}~{}\Delta\vec{D^{\prime}},$ (2.14)
@@
  • GRB: To obtain the likelihood for GRB, We compare both $\mu_{obs,GRBs}$ given in Equ. 2.4 with the theoretical $\mu_{th}$, defined as:
@@
    By combining the definitions in Equ. 2.3 and Equ. 2.5 we finally reproduce Equation 2.4. We then write the likelihood as:
@@
  • QSO: The likelihood used for QSOs used in our analysis is given as [see also 76, 77, 78, 79]:
    $ln(\mathcal{L}(QSO))=-\frac{1}{2}\sum_{i=1}^{N}\left[\frac{(y_{i}-\phi_{i})^{2}}{s^{2}_{i}}+\text{ln}(s^{2}_{i})\right].$ (2.17)
@@
Along with the 6 standard parameters in baseline $\Lambda$CDM model we have $1$ intrinsic parameter for Pantheon-Plus i.e. $M_{B}$, 3 intrinsic parameters for GRB i.e. $a,b,svGRB$ and 3 intrinsic parameters for QSO i.e. $g,be,svQ$. We have used publically available CLASS [80] code to calculated the background parameters. To implement the likelihoods we have used MontePython[81, 82] which uses Metropolis-Hastings algorithm to maximize the likelihoods and derive constraints on cosmological parameters as well as the intrinsic parameters. All of our runs were set up to reach a Gelman-Rubin convergence criterion of $R-1<10^{-2}$.
@@ 3.1 $M$ Tension
Figure 1: Contour plots of the values of $M$ and $H_{0}$ using the Planck+ Pantheon-Plus data and Pantheon-Plus combined with SH0ES are shown in dark green and light green, respectively.
@@ 3.3 Impact on GRB intrinsic parameters
Figure 3: The contour plots for the parameters $a$, $b$, $\sigma_{v}$, and $H_{0}$ for Planck combined with GRBs and for GRBs combined with Pantheon and with SH0ES data are shown in darker blue and lighter green, respectively.